# Exploring the topological sector optimization on quantum computers

### v2

@@ Abstract
Optimization problems are the core challenge in many fields of science and engineering, yet general and effective methods are scarce for searching optimal solutions. Quantum computing has been envisioned to help solve such problems, for example, the quantum annealing (QA) method based on adiabatic evolution has been extensively explored and successfully implemented on quantum simulators such as D-wave’s annealers and some Rydberg arrays. In this work, we investigate topological sector optimization (TSO) problem, which attracts particular interests in the quantum simulation and many-body physics community. We reveal that the topology induced by frustration in the optimization model is an intrinsic obstruction for QA and other traditional methods to approach the ground state. We demonstrate that the difficulties of TSO problem are not restricted to the gaplessness, but are also due to the topological nature which are often ignored for the analysis of optimization problems before. To solve TSO problems, we utilize quantum imaginary time evolution (QITE) with a possible realization on quantum computers, which exploits the property of quantum superposition to explore the full Hilbert space and can thus address optimization problems of topological nature. We report the performance of different quantum optimization algorithms on TSO problems and demonstrate that their capability to address optimization problems are distinct even when considering the quantum computational resources required for practical QITE implementations.
@@ I Introduction
Following technological advancements in manufacturing coupled qubit systems recently, QA can be implemented with superconducting flux qubits [17, 18, 19]. In the past decades, QA has been considered as a more powerful toolbox to solve the optimization problems than simulated annealing (SA) and other classical methods [5, 6, 20, 21, 22, 23, 24]. Furthermore, the Rydberg array simulator can also realize the Ising-like encoded Hamiltonian with high tunability and scalability [25, 26]. Despite of tremendous potentialities that QA possesses, it is yet an open question on which optimization problems can QA outperform other classical or quantum algorithms. Although the quantum fluctuations during the annealing process gives it more chance to jump out of an local minimum, the tangible limit of QA is hard to quantify. Especially when the ground states at two neighbor time moments are not connected smoothly, i.e., there is a phase transition in between, the validity of QA cannot always be guaranteed.
@@
Different from traditional optimization problems which focus on searching the ground state among many local minima in a glass system (see Fig. 1), the TSO problems further introduce several topological sectors in the energy landscape (see Fig. 1). Beside the challenges of local minima, the topological obstructions bring new difficulties for traditional quantum optimization algorithms because each topological sector has been protected and is robust to local quantum fluctuations. Generally speaking, a complex system with frustrations can emerge topological properties [33, 34, 35, 36, 37, 38, 39, 40] and its optimization problem naturally has the corresponding difficulties. Therefore, the TSO problems are important and urgent to be addressed universally.
@@ II.1 Quantum annealing and adiabaticity
$\frac{\max_{t\in[0,T]}\langle\chi_{n}(t)|\dot{\hat{H}}(t)|\chi_{m}(t)\rangle}{[\min_{t\in[0,T]}g(t)]^{2}}\ll 1$ (4)
should hold, where $\{|\chi_{n}(t)\rangle\}$ are the instantaneous orthogonal eigenvectors of $\hat{H}(t)$ and $\{E_{n}(t)\}$ are the corresponding eigenvalues. Therefore we must have $T\gg 1/g^{2}_{\text{min}}$, where $g_{\text{min}}\equiv\min_{t\in[0,T]}g(t)$. It is worth noting that the total time cost of the QA process $T$ should be larger than $1/g_{\text{min}}^{2}$ at least. We will find below that the time for QITE is naturally much less than QA according to its own mechanisms.
Based on the adiabatic approximation (4), the main obstacle of QA is accordingly the gaplessness. However, this is not the only factor that influences the efficacy of QA. If $\hat{H}_{0}$ has some different topological sectors, as we will show later, the performance of QA will dramatically decrease.
II.2 Sweeping quantum annealing
The basic idea of SQA is effectively opening a virtual edge through tuning a line-form annealing term [11]. The large transverse field acting on a global line can decrease the Ising-like interaction of the spins (qubits) on the line which is effective as open a virtual edge. Then the topological defect can be moved out through the open edge which effectively realizes a topological annihilation operator. In particular, we have to sweep this process in each meridian of the lattice, to guarantee all the topological defect can be touched. More details about this method applied in this work can be found in Appx. B.
II.3 Quantum imaginary time evolution
@@
The evolution of the instantaneous ground state in QA can be considered as searching the final ground state along a specific path in the Hilbert space following the adiabatic evolution. On the contrary, QITE is much more straightforward which can discard high energy states absolutely to explore the target state in a global way via the superposition property of a quantum system.
@@
Eq. (7) also indicates that the minimum evolutionary time $t$ to obtain a sufficiently pure ground state should scale with the reciprocal of the first excited state gap $g$, or $t\gg 1/g$, which offers a quadratic speed up compared to the real-time QA at least. It is worth noting that the $g$ here is the first excited gap in the $H_{t=T}$ while the $g_{\text{min}}$ in QA is the smallest gap during the whole evolution, thus the $g_{\text{min}}\leq g$. In most cases, we only need one (classical) solution of an optimization problem, and the gaplessness, implying the degeneracy of solution space, will not harm the application of QITE.
The infinite QITE operator ($t\to\infty$) is a projection operation onto the ground state(s) of $\hat{H}$ by decaying all other excited states. This method is free from problems like trap in a local minimum and lack of ergodicity. Moreover, there is no restriction about the specific form of the Hamiltonian such as short-range or long-range interactions, two-body or many-body interactions, and which manifold the spins form. Therefore, QITE is a more powerful tool for tackling TSO problem in principle. We also want to emphasize that the initial state of QITE is easy to be set as a superposition of all the final solutions in the Ising-like problem because all the solution is classical. What we need to do is just setting each qubit as $a|1\rangle+b|0\rangle$ with $ab\neq 0$.
II.4 Implementation of quantum imaginary time evolution
@@
The basic idea of VQITE is to prepare a trial state $|\phi(\vec{\theta}(t))\equiv|\phi(t)\rangle$ output by an circuit ansatz $V(\vec{\theta})$, a parameterized quantum circuit (PQC) on a quantum computer, i.e.
$|\phi(t)\rangle=V(\vec{\theta})|\bar{0}\rangle,$ (8)
where $\vec{\theta}=[\theta_{0},\theta_{1},\dots\theta_{p-1}]$ denotes $p$ variational parameters and $|\bar{0}\rangle$ is the inital state input into the PQC. To perform quantum imaginary time evolution, we require it to approximately satisfy the Wick-rotated Schrodinger equation under McLachlan’s variational principle [63]:
@@
$\mathbf{A}\vec{\dot{\theta}}=\mathbf{C}$ (10)
@@
can be evaluated on quantum computers via the Hadamard test. Eq. (11) will be futher solved with the help of a classical computer and the dynamics of $|\phi(t)\rangle$ is controlled by $\vec{\theta}$ as
@@ III Designing typical TSO problems
To test the efficiency of the aforementioned quantum optimization algorithms, several representative TSO problems are needed. The TSO model is expected to be simple enough in the Hamiltonian form, but sufficiently difficult to be optimized. In Fig. 2, we show the three characteristics of a difficult TSO problem: (i) The minimum energies of different topological sectors are nearly equal; (ii) The target topological sector containing the ground state occupies an exponentially small fraction of the Hilbert space thereupon is difficult to be found; (iii) Competing topological sectors have much larger Hilbert space to the benefit of quantum fluctuations. In other words, the search methods rely on quantum kinetic terms might render the competing topological sectors much more favored than the target one.
It is worth noting that the disordered terms are not necessary in the design, because our key point is to discuss the optimization between topological sectors. Moreover, the disordered problem (spin-glass) has been already widely studied in amount of previous works. With these characteristics in mind, we design the following two models and test the algorithms by numerical simulations.
@@ IV.2 Results
We perform real-time simulations of QA and SQA numerically, which can provide quantitative estimations for their practical performance on quantum annealers. We simulate the ideal QITE according to the definition Eq. (5) and on the other side, implement VQITE, Diag-VQITE and VQE using TensorCircuit quantum simulation framework [75]. We consider a $4\times 4$ lattice with $16$ qubits. Technical details of the simulations can be found in Appx. A, B and C.
@@
Notice that though the performance of VQITE should be better than Diag-VQITE as expected, it is even better than QITE according to Fig. 5 in this example. As a variational algorithm, the performance of VQITE depends closely on the choice of ansatz as introduced in the Sec. II.4. On the other hand, Diag-VQITE and VQE (see Fig. 5) can both find the ground state. The performance advantage of VQITE and the success of Diag-VQITE and VQE may come from the translation symmetry along the $x$-axis of the model, rendering the ground state easier to identify. Therefore, we consider a similar fully frustrated Ising model on square lattice, which breaks the translation symmetry in the next section.
@@ V.1 Model
Figure 6: (a) Lattice of $\hat{H}_{\text{sq}}$. The thick gray bonds correspond to the coupling $J$, the think black bonds correspond to the coupling $K$, and the blue think bonds correspond to the coupling $K_{p}$.
@@ V.2 Results
We use the same optimization schemes with $\hat{H}_{\text{tri}}$. As it is shown in Fig. 7 and 7, QA and SQA also fails in this TSO problem. However, different from $\hat{H}_{\text{tri}}$, $T=40$ is not even sufficient to reach the first excited state $E_{1}$ for QA. On the other hand, QITE needs $T\approx 4$ while VQITE needs $T<20$ to reach the ground state. In this example, both Diag-VQITE and VQE fail (see Fig. 5). Therefore the complete version of VQITE is necessary in this more difficult TSO problem. Besides, since the imaginary time evolution relates to the quantum natural gradient descent [64], its advantage over the stochastic gradient gradient by VQE is also showed in this model.
@@ V.3 The role of topology
We deliberately take $J=0.1$ to make sure that $\hat{H}_{\text{Ising}}$ shares the same gap with $\hat{H}_{\text{sq}}$ in the QA Hamiltonian Eq. (1). Consequently, the lower energies of $\hat{H}_{\text{eq}}$ and $\hat{H}_{\text{Ising}}$ computed by the exact diagonalization method show similar behaviors: both of the spectra have a closed gap at some point around $s=0.8$ (not the same one), and $E_{0}$ merges with $E_{1}$ after the gapless point, as it shown in Fig. 8.
Fig. 9 shows that when $T=10^{3}$, QA can find the ground state of $\hat{H}_{\text{Ising-2d}}$ exactly. Although there is a gapless point in the annealing path, the second order phase transition nature renders the critical dynamics following Kibble-Zurek mechanism, which secures the large fidelity with the ground state at the end [76, 77]. In $\hat{H}_{\text{sq}}$, the states before and after the gapless point are in different topological sectors, thus they are not connected smoothly hence it is hard for QA to remain correct.

### v3

@@ Abstract
Optimization problems are the core challenge in many fields of science and engineering, yet general and effective methods for finding optimal solutions remain scarce. Quantum computing has been envisioned to help solve such problems, with methods like quantum annealing (QA), grounded in adiabatic evolution, being extensively explored and successfully implemented on quantum simulators such as D-wave’s annealers and some Rydberg arrays. In this work, we investigate the topological sector optimization (TSO) problem, which attracts particular interests in the quantum simulation and many-body physics community. We reveal that the topology induced by frustration in the optimization model is an intrinsic obstruction for QA and other traditional methods to approach the ground state. We demonstrate that the difficulties of TSO problem are not restricted to the gaplessness, but are also due to the topological nature which are often ignored for the analysis of optimization problems before. To solve TSO problems, we utilize quantum imaginary time evolution (QITE) with a possible realization on quantum computers, which leverages the property of quantum superposition to explore the full Hilbert space and can thus address optimization problems of topological nature. We report the performance of different quantum optimization algorithms on TSO problems and demonstrate that their capability to address optimization problems are distinct even when considering the quantum computational resources required for practical QITE implementations.
@@ I Introduction
Following technological advancements in manufacturing coupled qubit systems recently, QA can be implemented with superconducting flux qubits [17, 18, 19]. In the past decades, QA has been considered as a more powerful toolbox to solve the optimization problems than simulated annealing (SA) and other classical methods [5, 6, 20, 21, 22, 23, 24]. Furthermore, the Rydberg array simulator can also realize the Ising-like encoded Hamiltonian with high tunability and scalability [25, 26]. Despite of tremendous potentialities that QA possesses, it is yet an open question on which optimization problems can QA outperform other classical or quantum algorithms. Although the quantum fluctuations during the annealing process gives it more chance to jump out of an local minimum, the tangible limit of QA is hard to quantify. Especially when the ground states at two adjacent time moments are not connected smoothly, i.e., there is a phase transition in between, the validity of QA cannot always be guaranteed.
@@
Different from traditional optimization problems which focus on searching the ground state among many local minima in a glass system (see Fig. 1), the TSO problems further introduce several topological sectors in the energy landscape (see Fig. 1). Beside the challenges of local minima, the topological obstructions bring new difficulties for traditional quantum optimization algorithms because each topological sector has been protected and is robust to local quantum fluctuations. Generally speaking, a complex system with frustrations can generate topological properties [33, 34, 35, 36, 37, 38, 39, 40] and its optimization problem naturally has the corresponding difficulties. Therefore, the TSO problems are important and urgent to be addressed universally.
@@ II.1 Quantum annealing and adiabaticity
Apparently, $\hat{H}_{1}$ only has a non-degenerate ground state
$|\Psi^{-}\rangle=\otimes_{j}|-\rangle_{j}=\otimes_{j}\bigg{(}\frac{|0\rangle_{j}-|1\rangle_{j}}{\sqrt{2}}\bigg{)},$ (4)
which is also a product state and easy to prepare practically.
@@
$\frac{\max_{t\in[0,T]}\langle\chi_{n}(t)|\dot{\hat{H}}(t)|\chi_{0}(t)\rangle\hbar}{[\min_{t\in[0,T]}g(t)]^{2}}\ll 1$ (5)
should hold for any $n\neq 0$, where $\{|\chi_{n}(t)\rangle\}$ are the instantaneous orthogonal eigenvectors of $\hat{H}(t)$ and $\{E_{n}(t)\}$ are the corresponding eigenvalues. Let $g_{\text{min}}\equiv\min_{t\in[0,T]}g(t)$ and suppose the numerator achieves its maximum at $t=t^{*}$. Then the variation in time at $t^{*}$ should satisfy $\Delta t\gg\hbar\langle\chi_{n}(t^{*})|\Delta\hat{H}|\chi_{0}(t^{*})\rangle/g_{\text{min}}^{2}\propto 1/g_{\text{min}}^{2}$. Therefore, to ensure the adiabacticity, the total evolution time must satisfy $T\gg\tau$, where $\tau\propto 1/g^{2}_{\min}$.
Based on the adiabatic approximation (5), the main obstacle of QA is accordingly the gaplessness. However, this is not the only factor that influences the efficacy of QA. As we will show later, if the Hilbert space of $\hat{H}_{0}$ comprises many topological sectors, the performance of QA will dramatically decrease. Relevant numerical evidence has also been presented in Ref. [28], where the authors numerically demonstrated the performance of QA, equivalently applied via digital quantum simulation, on the toric code model under external magnetic fields [48]. They found that the true topological sector of the toric code model cannot be achieved from a non-topological limit, which corresponds to the initial Hamiltonian $\hat{H}_{1}$.
II.2 Real-time simulations of quantum annealing
In this subsection, we introduce the method we used to perform our real-time simulations of QA on classical computers. For the QA Hamiltonian $\hat{H}(s)$, quantum mechanics requires that the immediate quantum state $|\Psi(t)\rangle$ and the state evolved after a small time interval $\Delta t$ satisfy the relation
$|\Psi(t+\Delta t)\rangle=e^{-i\hat{H}(s)\Delta t}|\Psi(t)\rangle.$ (6)
After this, we change $s$ by $\Delta s$ and perform the next evolution. Without loss of generality, we consider the case that $s$ linearly decreases with $t$. The total number of evolution steps is therefore $N=1/\Delta s$ and the total evolution time is $T=N\Delta t$. To ensure the adiabatic approximation, $T$ must be large enough or equivalently, $\Delta s/\Delta t$ is small enough when we use Eq. (6). For the two parameters $\Delta t$ and $\Delta s$, practically, to control the pace of QA, we can fix one of the them and change the other. In our simulations, we set $\Delta t=0.1$, and utilize different values of $\Delta s$ to adjust the total evolution time $T$. The initial state, as discussed before, is taken to be state (4).
For the first example and benchmark, we test it on a 16-qubit Ising chain under periodic boundary condition, whose Hamiltonian is written as
As it shown in Fig. (2), when $T=40$, the process is non-adiabatic with great fluctuations when $s\to 1$. It turns out that $\Delta s=10^{-4}$, i.e. $T=10^{3}$, is sufficient for this model to reach its ground state $E_{0}$ via QA.
Figure 2: Real-time simulation of QA for a 16-qubit Ising chain under periodic boundary condition with different evolution time. Here we take $\hat{H}_{0}\equiv\hat{H}_{\text{Ising-1d}}$. The black dashed line denotes the exact ground state energy of the model.
II.3 Sweeping quantum annealing
As we will introduce in Sec. IV and V, different topological sectors are separated and marked by some global topological defects, which in spin systems are usually some domain walls. If we possess prior knowledge of the direction of the defects, we can polarize the spins on a line (called virtual edge) parallel to the defect on the lattice using strong transverse magnetic fields. These fields can disrupt the defect when it overlaps with the virtual edge, enabling the state to escape the confines of that particular topological sector. Since practically we cannot assure the defect can overlap the virtual edge we choose, we adopt the strategy that sweeping all parallel lines and making them to be the virtual edge sequentially in the process of a vanilla QA. This method is called the sweeping quantum annealing (SQA) [11], effectively introducing some topological annihilation operators with a high likelihood of eliminating the defect.
However, since the prior knowledge cannot always be assumed, and furthermore, if the true ground state lies within a topological sector that has many topological defects, the creation of virtual edges would be counterproductive. We also explore this method in our work, employing the same initial state as in QA. More numerical details about the real-time simulation of SQA can be found in Appx. A.
II.4 Quantum imaginary time evolution
@@
The evolution of the instantaneous ground state in QA can be considered as searching the final ground state along a specific path in the Hilbert space following the adiabatic evolution. On the contrary, QITE offers a more direct approach, capable of discarding high energy states absolutely over time, thereby enabling the exploration of the target state in a global manner through the superposition property inherent in quantum systems.
@@
Eq. (10) indicates that the minimum evolutionary time $t$ to obtain a sufficiently pure ground state should scale with the reciprocal of the first excited state gap $g$, or $t\gg 1/g$, which offers a quadratic speed up compared to the real-time QA at least. It is worth noting that the $g$ here is the first excited gap in the $H_{t=T}$ while the $g_{\text{min}}$ in QA is the smallest gap during the whole evolution, thus the $g_{\text{min}}\leq g$. In most cases, we only need one (classical) solution of an optimization problem, and the gaplessness, implying the degeneracy of solution space, will not harm the application of QITE.
Moreover, Eq. (10) provides a mathematical representation of the infinite QITE operator ($t\to\infty$), illustrating its role as a projection operation onto the ground state(s) of $\hat{H}$ by diminishing all other excited states. Therefore, this method is free from problems such as being trapped in a local minimum or lacking ergodicity, since all eigenstates, including our desired solutions, are equally addressed simultaneously. Moreover, this projection imposes no constraints on the specific characteristics of the Hamiltonian, be it short-range or long-range interactions, involving two-body or many-body interactions, or the lattice manifold [49, 50]. Thus, QITE is a more powerful tool for tackling TSO problem theoretically. Since we cannot assume prior knowledge of the ground state, it becomes imperative to consider a superposition of all possible classical states within the Hilbert space as the initial state. In our work, we take
$|\Psi^{+}\rangle=\otimes_{j}|+\rangle_{j}=\otimes_{j}\bigg{(}\frac{|0\rangle_{j}+|1\rangle_{j}}{\sqrt{2}}\bigg{)},$ (11)
for convenience.
II.5 Implementation of quantum imaginary time evolution
@@
To perform VQITE, we first prepare a trial state
$|\phi(\vec{\theta}(t))\equiv|\phi(t)\rangle=V(\vec{\theta})|\bar{0}\rangle,$ (12)
where $V(\vec{\theta})$ is some circuit ansatz, a parameterized quantum circuit (PQC) on a quantum computer; $\vec{\theta}=[\theta_{0},\theta_{1},\dots\theta_{p-1}]$ denotes $p$ variational parameters and $|\bar{0}\rangle$ is the inital state input into the PQC. Thereafter, we require it to approximately satisfy the Wick-rotated Schrodinger equation under McLachlan’s variational principle [65]:
@@
$\mathbf{A}\dot{\vec{\theta}}=\mathbf{C}$ (14)
@@
The $A_{jk}$ and $C_{j}$ in Eq. (16) can be evaluated on a quantum computer via the Hadamard test, then Eq. (15) will be futher solved with the help of a classical computer and the dynamics of $|\phi(t)\rangle$ is controlled by $\vec{\theta}$ as
@@ III Designing typical TSO problems
To test the efficiency of the aforementioned quantum optimization algorithms, several representative TSO problems are needed. The TSO model is expected to be simple enough in the Hamiltonian form, but sufficiently difficult to be optimized. Models that are amenable to numerical simulation on classical computers, expected to involve around 16 qubits or more to make topological effects manifest, are rare. We have identified two suitable examples for our study, which will be introduced in Sec. IV and Sec. V. Many other models would necessitate greater computational resources, posing challenges for simulating variational algorithms that require substantial memory to store the variational parameters in addition to the qubits.
Moreover, these two models represent a challenging scenario even within the class of TSO problems. If QITE proves effective in such a scenario where QA struggles, it would underscore the advantage of QITE. In Fig. 3, we show the three characteristics of this kind of TSO problems: (i) The minimum energies of different topological sectors are nearly equal; (ii) The target topological sector containing the ground state occupies an exponentially small fraction of the Hilbert space thereupon is difficult to be found; (iii) Competing topological sectors have much larger Hilbert space to the benefit of quantum fluctuations. In other words, utilizing search methods based on quantum kinetic terms could potentially give rise to a situation where the competing topological sectors become significantly more preferred compared to the desired target sector.
There are three reasons for the above criteria. 1) The case we designed for tests should be a difficult task in the topological optimization problem. Therefore, it is natural to set the degree of freedom of the target sector as small as possible. It can avoid the case that the target can be found easily via random searching. 2) The model needs being simple in understanding, so we do not consider the spin glass problem but focused on the topological optimization itself only. This makes the problem more focused, eliminating the effects of other variables, and the well-known ground state of the model is easy for benchmark. 3) This hardcore mode has been discussed carefully in Ref.[11] by different QA methods, thus it is convenient for comparing the advantages of the QITE with QAs. With these characteristics in mind, we design the following two models and test the algorithms by numerical simulations.
@@ IV.2 Results
We perform real-time simulations of QA and SQA numerically, which can provide quantitative estimations for their practical performance on quantum annealers. We simulate the ideal QITE according to the definition Eq. (8) and on the other side, implement VQITE, Diag-VQITE and VQE using TensorCircuit quantum simulation framework [77]. We consider a $4\times 4$ lattice with $16$ qubits. More technical details of the simulations can be found in Appx. A and B. In addition, all codes for numerical simulations and results shown in this paper are open-source  [^1] https://github.com/msquare1998/SourceCodes-arxiv2310.04291.
@@
Notice that though the performance of VQITE should be better than Diag-VQITE as expected, it is even better than QITE according to Fig. 6 in this example. As a variational algorithm, the performance of VQITE depends closely on the choice of ansatz as introduced in the Sec. II.5. On the other hand, Diag-VQITE and VQE (see Fig. 6) can both find the ground state. The performance advantage of VQITE and the success of Diag-VQITE and VQE may come from the translation symmetry along the $x$-axis of the model, rendering the ground state easier to identify. Therefore, we consider a similar fully frustrated Ising model on square lattice, which breaks the translation symmetry in the next section.
@@ V.1 Model
Figure 7: (a) Lattice of $\hat{H}_{\text{sq}}$. The thick gray bonds correspond to the coupling $J$, the think black bonds correspond to the coupling $K$, and the blue think bonds correspond to the coupling $K_{p}$. (b) the dimer model on the dual lattice which is depicted in pink color with the green links denoting the dimers.
@@ V.2 Results
We use the same optimization schemes with $\hat{H}_{\text{tri}}$. As it is shown in Fig. 8 and 8, QA and SQA also fails in this TSO problem. However, different from $\hat{H}_{\text{tri}}$, $T=40$ is not even sufficient to reach the first excited state $E_{1}$ for QA. On the other hand, QITE needs $T\approx 4$ while VQITE needs $T<20$ to reach the ground state. In this example, both Diag-VQITE and VQE fail (see Fig. 6). Therefore the complete version of VQITE is necessary in this more difficult TSO problem. Besides, since the imaginary time evolution relates to the quantum natural gradient descent [66], its advantage over the stochastic gradient gradient by VQE is also showed in this model.
@@ V.3 The role of topology
We take $J=0.1$ to make sure that $\hat{H}_{\text{Ising}}$ shares the same gap with $\hat{H}_{\text{sq}}$ in the QA Hamiltonian Eq. (1). Consequently, the lower energies of $\hat{H}_{\text{eq}}$ and $\hat{H}_{\text{Ising}}$ computed by the exact diagonalization method show similar behaviors: both of the spectra have a closed gap at some point around $s=0.8$ (not the same one), and $E_{0}$ merges with $E_{1}$ after the gapless point, as it shown in Fig. 9.
Fig. 10 shows that when $T=10^{3}$, QA can find the ground state of $\hat{H}_{\text{Ising-2d}}$ exactly. Although there is a point that gap closes in the annealing path, the second order phase transition nature renders the critical dynamics following Kibble-Zurek mechanism, which secures the large fidelity with the ground state at the end [79, 80]. In $\hat{H}_{\text{sq}}$, the states before and after the gapless point are in different topological sectors, thus they are not connected smoothly hence it is hard for QA to remain correct.
QA with the Hamiltonian in Eq.(1) leverages quantum fluctuations induced by magnetic fields along the $x$ direction to seek the true ground state. Since the magnetic fields are local (i.e. each $\hat{X}_{i}$ only acts on a single site $i$), they have little chance to induce a collective fluctuation that aligns precisely with the global topological defect to be bypassed for an instaneous state. As the system size increases, the length of the defect grows, decreasing the likelihood of such collective fluctuation. We can also understand this point through general topological phases [81, 82], in which the topological sectors are robust to local magnetic perturbations.